OpenStudy (anonymous):

Could someone explain how to reduce 8/sqr2

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

rationalize the denominator by mutliplying the entire fraction by sqr2/sqr2 you can do this because sqr2/sqr2 = 1

OpenStudy (tkhunny):

\(\dfrac{8}{\sqrt{2}}\) is already simplified. However, you can change it around a little to make someone happy. \(\dfrac{8}{\sqrt{2}} = \dfrac{8}{\sqrt{2}}\cdot\dfrac{\sqrt{2}}{\sqrt{2}} = \dfrac{8\sqrt{2}}{2} = \dfrac{8}{2}\sqrt{2} = 4\sqrt{2}\)